OpenStudy (anonymous):

What is the factored form of the expression? 3x2 + 8x – 16 A. (3x + 4)(x – 4) B. (3x + 4)(x + 4) C. (3x – 4)(x – 4) D. (3x – 4)(x + 4)

OpenStudy (anonymous):

@RadEn help

OpenStudy (anonymous):

i got A

OpenStudy (anonymous):

not sure

OpenStudy (raden):

3x^2 = 3x * x 3x*4 + (-4)x = 8x -4*4 = -16 so, it is D (3x – 4)(x + 4)

OpenStudy (anonymous):

ok thanks

OpenStudy (raden):

you're welcome